Some hydrogen gas a volume of 250cm3 at 20°c if the pressure is kept constant at what temperature will it's volume be 100cm3?





1
Expert's answer
2021-12-21T09:20:01-0500

"T_1= 273+20=293k"


"\frac{V_1}{T_1}=
\frac{V_2}{T_2}"


"T_2=
\frac{V_2×T_1}{V_1}"


"T_1
=
\frac{293×100}{250}=117.2k=-155.8°c"